Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

2021-02-22

Observers:  Luke Harvey, Simon Prentice
Support: Maxime Deckers, Mark Magee
TAT: Steve Schulze, Jacob Teffs, Santiago Gonzalez

EFOSC Calibrations: bias, calib, UBVRI flats, gri flats


Weather:
Start: humidity 30, wind 8.7, seeing 1.4
Middle: humidity 20, wind 12.6, seeing 1.2
End: humidity 6, wind 12.8, seeing -

Comments: 
Clouds from 01:15 - 02:50.
Clouds from 03:30 - 05:40.
TO P2 not updating, unable to add OBs.
Wind caused pointing restrictions from 5:40.
The seeing monitor was not operational for the latter half of the night.


Started with EFOSC

 UT Target Type Set up
 Slit Exposure Airmass Seeing Notes
00:27SN2020abjxFollow-upGr111.02700s1.81.2Some cloud interference with guide star visibility. No trace seen in the spectrum.
02:56L745aStandardGr11, Gr13, Gr161.0 1.580s1.00.9
03:09AT2021ckjFollow-up (photometry)BVRI-300s1.11.1Aborted due to clouds
05:44AT2021ckjFollow-up (photometry)BVRI-300s1.3 Restarted after reopening the dome.
06:43SN2021dovFollow-upGr131.02700s1.6  
07:35SN2021dbgFollow-upGr111.01500s1.8  
08:05SN2020acatFollow-up Gr11, Gr161.01500s1.2  
Comments